OpenStudy (anonymous):

(1-squareroot 5)(1+squareroot 5)

OpenStudy (anonymous):

(a-b)(a+b) = a^2 - b^2 ( sqrt(5) )^2 = 5 1 ^2 = 1

OpenStudy (anonymous):

Im confused so the answer is 5 squareroot 1?

OpenStudy (anonymous):

no I told you what a^2 is and what b^2 is, you just have to subtract them to get the final answer

OpenStudy (anonymous):

So sqaureroot 4?

OpenStudy (anonymous):

are u guessing here? why sqrt(4) ? 1-5 = ?

OpenStudy (anonymous):

Sorry I was thinking 5-1 so the answer is squareroot -4

OpenStudy (anonymous):

there is no sqrt its just -4

OpenStudy (anonymous):

Oh wow thats exactly what I had at firts I second guessed myself.......Thanks for helping me! :)

OpenStudy (anonymous):

Can I ask one more question?

OpenStudy (anonymous):

pls post it as a new question and not here under this one

OpenStudy (anonymous):

Okay